# Mapping 1,000+ Language Models via the Log-Likelihood Vector

### v1

@@ Appendix I Table of Nearest Neighbor Models
     Pearson’s $r$ 0.969$\pm$0.002 0.943$\pm$0.004 0.959$\pm$0.003 0.954$\pm$0.001 0.961$\pm$0.002 0.930$\pm$0.002 0.973$\pm$0.001 0.993$\pm$0.001
     Spearman’s $\rho$ 0.975$\pm$0.001 0.971$\pm$0.002 0.966$\pm$0.003 0.933$\pm$0.001 0.971$\pm$0.002 0.892$\pm$0.001 0.976$\pm$0.001 0.989$\pm$0.001
Table 6: This table presents the mean and standard deviation of correlation coefficients between predicted and actual benchmark scores. These coefficients were computed using ridge regression across five different data splits. Results for predicting 6-TaskMean and the mean log-likelihood are also included.
@@
For the setting where the target variable $\mathbf{f}$ was replaced with the mean log-likelihood (¯$(\bar{\ell}_{1},\ldots,\bar{\ell}_{K})\in\mathbb{R}^{K}$, we searched for $\alpha$ within $\{10^{-4},\ldots,10^{4}\}$ and did not apply clipping as a post-processing step.
@@
To account for randomness, we ran five different data splits when predicting each benchmark score. The final predicted score was the average of these five runs. For each split, we computed the correlation coefficients between the predicted scores and the actual benchmark scores. Table 6 presents their mean and standard deviation, and shows a similar trend as Table 2.
@@
Table LABEL:tab:model_table lists the 1,018 models used in this study, sorted alphabetically by their names. The BibTeX entries cited for each model were determined through the following procedure.
@@
1 aaditya/Llama3-OpenBioLLM-8B Singhal et al. (2022); Nori et al. (2023); Singhal et al. (2023); Pal and Sankarasubbu (2024a, b); Rafailov et al. (2024) llama-3 8B 2024-04-20 9308 -590.14 54.06
@@
90 berkeley-nest/Starling-LM-7B-alpha Zhu et al. (2023b, a) mistral 7B 2023-11-25 18366 -571.63 67.05
@@
95 bigcode/starcoder2-3b Beltagy et al. (2020); Bavarian et al. (2022); Dao et al. (2022); Ainslie et al. (2023); Lozhkov et al. (2024) starcoder2 3B 2023-11-29 445316 -620.70 39.25
96 bigcode/starcoder2-7b Beltagy et al. (2020); Bavarian et al. (2022); Dao et al. (2022); Ainslie et al. (2023); Lozhkov et al. (2024) starcoder2 7B 2024-02-20 11834 -584.03 42.95
@@
100 bigscience/bloom-1b1 Shoeybi et al. (2020); Press et al. (2022); Dettmers et al. (2022) bloom 1B 2022-05-19 11054 -687.97 32.47
101 bigscience/bloom-1b7 Shoeybi et al. (2020); Press et al. (2022); Dettmers et al. (2022) bloom 1B 2022-05-19 40840 -657.05 33.98
102 bigscience/bloom-3b Shoeybi et al. (2020); Press et al. (2022); Dettmers et al. (2022) bloom 3B 2022-05-19 13914 -633.79 36.07
103 bigscience/bloom-7b1 Shoeybi et al. (2020); Press et al. (2022); Dettmers et al. (2022) bloom 7B 2022-05-19 21428 -604.20 39.18
@@
242 dreamgen/WizardLM-2-7B Luo et al. (2023); Xu et al. (2023a); Luo et al. (2025) mistral 7B 2024-04-16 2533 -641.93 63.75
@@
253 EleutherAI/polyglot-ko-12.8b Kim et al. (2022); Ko et al. (2023); Su et al. (2023a) gpt_neox 13B 2022-10-14 2941 -967.65 33.33
@@
263 elyza/ELYZA-japanese-Llama-2-13b-instruct Touvron et al. (2023b); Sasaki et al. (2023b) llama-2 13B 2023-12-25 1710 -594.98 54.72
264 elyza/ELYZA-japanese-Llama-2-13b Touvron et al. (2023b); Sasaki et al. (2023b) llama-2 13B 2023-12-25 1200 -587.25 56.14
@@
308 GeneZC/MiniChat-1.5-3B Jain et al. (2023); Zhang et al. (2024a); Rafailov et al. (2024) llama 3B 2023-11-26 1601 -599.51 50.23
309 GeneZC/MiniChat-2-3B Jain et al. (2023); Zhang et al. (2024a); Rafailov et al. (2024) llama 3B 2023-12-27 4167 -607.10 51.49
@@
319 google/gemma-1.1-7b-it Joshi et al. (2017); Zhao et al. (2018); Mihaylov et al. (2018); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Bisk et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Zellers et al. (2019); Austin et al. (2021); Cobbe et al. (2021); Hendrycks et al. (2021a); Chen et al. (2021); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-03-26 19835 -1355.60 60.09
@@
320 google/gemma-2b-it Joshi et al. (2017); Mihaylov et al. (2018); Rudinger et al. (2018); Zhao et al. (2018); Sap et al. (2019); Bisk et al. (2019); Zellers et al. (2019); Clark et al. (2019); Talmor et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Cobbe et al. (2021); Hendrycks et al. (2021a); Dhamala et al. (2021); Chen et al. (2021); Austin et al. (2021); Hartvigsen et al. (2022); Parrish et al. (2022); Lin et al. (2022a); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 2B 2024-02-08 103851 -886.70 42.75
@@
321 google/gemma-2b Joshi et al. (2017); Mihaylov et al. (2018); Rudinger et al. (2018); Zhao et al. (2018); Sap et al. (2019); Bisk et al. (2019); Zellers et al. (2019); Clark et al. (2019); Talmor et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Cobbe et al. (2021); Hendrycks et al. (2021a); Dhamala et al. (2021); Chen et al. (2021); Austin et al. (2021); Hartvigsen et al. (2022); Parrish et al. (2022); Lin et al. (2022a); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 2B 2024-02-08 256798 -581.35 46.51
322 google/gemma-7b-it Joshi et al. (2017); Mihaylov et al. (2018); Rudinger et al. (2018); Zhao et al. (2018); Sap et al. (2019); Bisk et al. (2019); Zellers et al. (2019); Clark et al. (2019); Talmor et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Cobbe et al. (2021); Hendrycks et al. (2021a); Dhamala et al. (2021); Chen et al. (2021); Austin et al. (2021); Hartvigsen et al. (2022); Parrish et al. (2022); Lin et al. (2022a); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-02-13 66776 -1327.43 53.56
@@
323 google/gemma-7b Joshi et al. (2017); Mihaylov et al. (2018); Rudinger et al. (2018); Zhao et al. (2018); Sap et al. (2019); Bisk et al. (2019); Zellers et al. (2019); Clark et al. (2019); Talmor et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Cobbe et al. (2021); Hendrycks et al. (2021a); Dhamala et al. (2021); Chen et al. (2021); Austin et al. (2021); Hartvigsen et al. (2022); Parrish et al. (2022); Lin et al. (2022a); Srivastava et al. (2023); Zhong et al. (2023); Dettmers et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-02-08 72047 -539.12 63.75
@@
324 google/recurrentgemma-2b-it Joshi et al. (2017); Mihaylov et al. (2018); Rudinger et al. (2018); Zhao et al. (2018); Sap et al. (2019); Bisk et al. (2019); Zellers et al. (2019); Clark et al. (2019); Talmor et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Cobbe et al. (2021); Hendrycks et al. (2021a); Dhamala et al. (2021); Chen et al. (2021); Hendrycks et al. (2021b); Austin et al. (2021); Hartvigsen et al. (2022); Parrish et al. (2022); Lin et al. (2022a); Srivastava et al. (2023); Zhong et al. (2023); De et al. (2024); Griffin Team et al. (2024) recurrent_gemma 2B 2024-04-08 4046 -1033.09 40.86
@@
325 gradientai/Llama-3-8B-Instruct-Gradient-1048k Peng et al. (2023b); Ding et al. (2023); Pekelis et al. (2024); AI@Meta (2024); Liu et al. (2025) llama-3 8B 2024-04-29 6854 -564.98 59.84
326 gradientai/Llama-3-8B-Instruct-262k Peng et al. (2023b); Ding et al. (2023); Pekelis et al. (2024); AI@Meta (2024); Liu et al. (2025) llama-3 8B 2024-04-25 12536 -555.17 60.26
@@
366 HuggingFaceH4/zephyr-7b-alpha Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistral 7B 2023-10-09 12911 -559.73 59.50
367 HuggingFaceH4/zephyr-7b-beta Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistral 7B 2023-10-26 294019 -570.40 59.08
@@
375 hywu/Camelidae-8x13B Houlsby et al. (2019); Komatsuzaki et al. (2023); Dettmers et al. (2023); Wu et al. (2024) camelidae 13B 2024-01-10 1886 -535.46 59.40
376 hywu/Camelidae-8x7B Houlsby et al. (2019); Komatsuzaki et al. (2023); Dettmers et al. (2023); Wu et al. (2024) camelidae 7B 2024-01-10 1901 -565.23 54.47
@@
387 IDEA-CCNL/Ziya-LLaMA-13B-Pretrain-v1 IDEA-CCNL (2021); Zhang et al. (2022a); Yang et al. (2022) llama-1 13B 2023-06-01 1171 -1397.30 29.96
388 IDEA-CCNL/Ziya-LLaMA-13B-v1 IDEA-CCNL (2021); Zhang et al. (2022a); Yang et al. (2022) llama-1 13B 2023-05-16 1244 -1397.34 29.82
@@
423 jphme/orca_mini_v2_ger_7b Conover et al. (2023); Mathur (2023a); Su et al. (2023b); Harries (2023); Taori et al. (2023); Touvron et al. (2023a); Xu et al. (2023a) llama-1 7B 2023-07-04 1181 -603.44 47.65
@@
566 meta-math/MetaMath-Llemma-7B Azerbayev et al. (2024); Yu et al. (2024b) llama 7B 2023-11-19 1322 -612.21 53.19
@@
601 mosaicml/mpt-7b-chat Henry et al. (2020); Dao et al. (2022); Press et al. (2022); MosaicML NLP Team (2023b) mpt 7B 2023-05-04 88069 -647.13 45.39
602 mosaicml/mpt-7b-instruct Henry et al. (2020); Dao et al. (2022); Press et al. (2022); MosaicML NLP Team (2023b) mpt 7B 2023-05-05 8599 -569.62 44.83
603 mosaicml/mpt-7b-storywriter Dao et al. (2022); Press et al. (2022); Chen et al. (2023a); MosaicML NLP Team (2023b) mpt 7B 2023-05-04 1920 -650.05 39.31
604 mosaicml/mpt-7b-8k-chat Henry et al. (2020); Dao et al. (2022); Press et al. (2022); MosaicML NLP Team (2023a) mpt 7B 2023-06-22 1304 -586.09 47.78
605 mosaicml/mpt-7b-8k Henry et al. (2020); Shoeybi et al. (2020); Press et al. (2022); Dao et al. (2022); Chen et al. (2023a); MosaicML NLP Team (2023b); Touvron et al. (2023a) mpt 7B 2023-06-30 2106 -546.57 47.24
606 mosaicml/mpt-7b Henry et al. (2020); Shoeybi et al. (2020); Press et al. (2022); Dao et al. (2022); Chen et al. (2023a); MosaicML NLP Team (2023b); Touvron et al. (2023a) mpt 7B 2023-05-05 31480 -548.53 44.28
@@
651 openaccess-ai-collective/jackalope-7b Mukherjee et al. (2023); Lian et al. (2023b); Longpre et al. (2023) mistral 7B 2023-10-07 1198 -550.33 61.16
@@
687 OpenModels4all/gemma-1.1-7b-it Joshi et al. (2017); Zhao et al. (2018); Mihaylov et al. (2018); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Bisk et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Zellers et al. (2019); Austin et al. (2021); Cobbe et al. (2021); Hendrycks et al. (2021a); Chen et al. (2021); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-04-06 5128 -1355.60 59.78
@@
692 Open-Orca/LlongOrca-13B-16k Mukherjee et al. (2023); Wang et al. (2023a); Dale et al. (2023); Longpre et al. (2023); Touvron et al. (2023b) llama-2 13B 2023-08-16 1217 -551.07 56.59
693 Open-Orca/LlongOrca-7B-16k Mukherjee et al. (2023); Wang et al. (2023a); Longpre et al. (2023); Touvron et al. (2023b); Lian et al. (2023c) llama-2 7B 2023-08-05 1223 -583.06 53.02
@@
695 Open-Orca/Mistral-7B-SlimOrca Mukherjee et al. (2023); Lian et al. (2023f, e); Longpre et al. (2023) mistral 7B 2023-10-08 3696 -569.42 60.37
696 Open-Orca/OpenOrcaxOpenChat-Preview2-13B Mukherjee et al. (2023); Wang et al. (2023a, b); Longpre et al. (2023); Touvron et al. (2023b) llama-2 13B 2023-07-31 2172 -547.85 56.70
697 Open-Orca/OpenOrca-Platypus2-13B Hu et al. (2022); Mukherjee et al. (2023); Wang et al. (2023a, b); Lee et al. (2023); Longpre et al. (2023); Touvron et al. (2023b); Lee et al. (2024) llama 13B 2023-08-11 6853 -557.72 57.28
698 Open-Orca/OpenOrca-Preview1-13B Mukherjee et al. (2023); Lian et al. (2023a); Longpre et al. (2023); Touvron et al. (2023a) llama-1 13B 2023-07-12 1240 -679.87 51.38
@@
742 rinna/youri-7b Andonian et al. (2021); Touvron et al. (2023b); Zhao et al. (2023a); Sawada et al. (2024) llama-2 7B 2023-10-30 2512 -545.03 47.11

### v2

@@ Appendix I Table of Nearest Neighbor Models
     Pearson’s $r$ $0.941\pm 0.002$ $0.904\pm 0.005$ $0.924\pm 0.006$ $0.889\pm 0.013$ $0.934\pm 0.005$ $0.864\pm 0.019$ $0.947\pm 0.005$ $0.987\pm 0.006$
     Spearman’s $\rho$ $0.944\pm 0.005$ $0.951\pm 0.003$ $0.926\pm 0.005$ $0.875\pm 0.003$ $0.943\pm 0.009$ $0.841\pm 0.013$ $0.956\pm 0.004$ $0.971\pm 0.006$
Table 7: Group 5-fold based on model types using $\bm{Q}$: Mean and standard deviation of the correlation coefficients between predicted and actual benchmark scores. Coefficients were obtained by ridge regression under five data splits based on model types. Results for predicting 6-TaskMean and the mean log-likelihood are also included.
     Pearson’s $r$ $0.968\pm 0.001$ $0.939\pm 0.004$ $0.960\pm 0.002$ $0.952\pm 0.001$ $0.960\pm 0.003$ $0.931\pm 0.001$ $0.973\pm 0.001$ $0.994\pm 0.001$
     Spearman’s $\rho$ $0.972\pm 0.001$ $0.970\pm 0.002$ $0.966\pm 0.001$ $0.929\pm 0.001$ $0.969\pm 0.001$ $0.891\pm 0.004$ $0.976\pm 0.001$ $0.990\pm 0.000$
Table 8: Random 5-fold using $\bm{Q}$: Mean and standard deviation of the correlation coefficients, obtained as in Table 7, but using five random data splits instead of splits based on model types.
     Pearson’s $r$ $0.939\pm 0.002$ $0.903\pm 0.007$ $0.923\pm 0.006$ $0.890\pm 0.013$ $0.935\pm 0.006$ $0.863\pm 0.018$ $0.945\pm 0.005$ $1.000\pm 0.000$
     Spearman’s $\rho$ $0.943\pm 0.005$ $0.952\pm 0.003$ $0.925\pm 0.005$ $0.875\pm 0.003$ $0.944\pm 0.008$ $0.841\pm 0.013$ $0.954\pm 0.004$ $1.000\pm 0.000$
Table 9: Group 5-fold based on model types using $\bm{L}$: Mean and standard deviation of the correlation coefficients, obtained as in Table 7, but using the matrix $\bm{L}$ instead of the matrix $\bm{Q}$ in the ridge regression of (5) in Section 5.
@@
For the setting where the target variable $\bm{v}\in\mathbb{R}^{K}$ was replaced with the mean log-likelihood (¯$(\bar{\ell}_{1},\ldots,\bar{\ell}_{K})\in\mathbb{R}^{K}$, we searched for $\alpha$ within $\{10^{-4},\ldots,10^{4}\}$ and did not apply clipping as a post-processing step.
@@
To account for randomness, we ran five different data splits based on model types when predicting each benchmark score. As explained in Section 5, the final predicted score was the average of these five runs. Additionaly, for each split, we computed the correlation coefficients between the predicted scores and the actual benchmark scores. Table 7 presents their mean and standard deviation, and shows a similar trend as Table 2.
K.3 Results from different settings
To investigate potential data leakage due to model types, Table 8 shows the correlation coefficients using five random data splits. These results demonstrate higher correlation coefficients across all tasks compared to those obtained using splits based on model types (Table 7), suggesting that randomly splitting models may unintentionally simplify the prediction task due to leakage from model types.
As shown in (5) in Section 5, we performed ridge regression using the double-centered log-likelihood matrix $\bm{Q}$, derived from the log-likelihood matrix $\bm{L}$. To assess the effect of replacing $\bm{Q}$ with $\bm{L}$, we repeated the analysis using $\bm{L}$ and report the resulting means and standard deviations of the correlation coefficients in Table 9. The results obtained with $\bm{L}$ (Table 9) show nearly the same trend as those with $\bm{Q}$ (Table 7). Note, however, that the mean of each row of $\bm{L}$ equals the mean log-likelihood itself, so this specific target can be reproduced exactly. Consequently, the mean correlation coefficients are $1.000$ for both Pearson’s $r$ and Spearman’s $\rho$.
@@
Table LABEL:tab:model_list lists the 1,018 models used in this study, sorted alphabetically by their names. The BibTeX entries cited for each model were determined through the following procedure.
@@
1 aaditya/Llama3-OpenBioLLM-8B Singhal et al. (2022, 2023); Nori et al. (2023); Rafailov et al. (2024); Pal and Sankarasubbu (2024b, a) llama-3 8B 2024-04-20 9308 -590.14 54.06
@@
90 berkeley-nest/Starling-LM-7B-alpha Zhu et al. (2023a, b) mistral 7B 2023-11-25 18366 -571.63 67.05
@@
95 bigcode/starcoder2-3b Beltagy et al. (2020); Dao et al. (2022); Bavarian et al. (2022); Ainslie et al. (2023); Lozhkov et al. (2024) starcoder2 3B 2023-11-29 445316 -620.70 39.25
96 bigcode/starcoder2-7b Beltagy et al. (2020); Dao et al. (2022); Bavarian et al. (2022); Ainslie et al. (2023); Lozhkov et al. (2024) starcoder2 7B 2024-02-20 11834 -584.03 42.95
@@
100 bigscience/bloom-1b1 Shoeybi et al. (2020); Dettmers et al. (2022); Press et al. (2022) bloom 1B 2022-05-19 11054 -687.97 32.47
101 bigscience/bloom-1b7 Shoeybi et al. (2020); Dettmers et al. (2022); Press et al. (2022) bloom 1B 2022-05-19 40840 -657.05 33.98
102 bigscience/bloom-3b Shoeybi et al. (2020); Dettmers et al. (2022); Press et al. (2022) bloom 3B 2022-05-19 13914 -633.79 36.07
103 bigscience/bloom-7b1 Shoeybi et al. (2020); Dettmers et al. (2022); Press et al. (2022) bloom 7B 2022-05-19 21428 -604.20 39.18
@@
242 dreamgen/WizardLM-2-7B Xu et al. (2023a); Luo et al. (2023, 2025) mistral 7B 2024-04-16 2533 -641.93 63.75
@@
253 EleutherAI/polyglot-ko-12.8b Kim et al. (2022); Su et al. (2023a); Ko et al. (2023) gpt_neox 13B 2022-10-14 2941 -967.65 33.33
@@
263 elyza/ELYZA-japanese-Llama-2-13b Sasaki et al. (2023b); Touvron et al. (2023b) llama-2 13B 2023-12-25 1200 -587.25 56.14
264 elyza/ELYZA-japanese-Llama-2-13b-instruct Sasaki et al. (2023b); Touvron et al. (2023b) llama-2 13B 2023-12-25 1710 -594.98 54.72
@@
308 GeneZC/MiniChat-1.5-3B Jain et al. (2023); Rafailov et al. (2024); Zhang et al. (2024a) llama 3B 2023-11-26 1601 -599.51 50.23
309 GeneZC/MiniChat-2-3B Jain et al. (2023); Rafailov et al. (2024); Zhang et al. (2024a) llama 3B 2023-12-27 4167 -607.10 51.49
@@
319 google/gemma-1.1-7b-it Joshi et al. (2017); Zhao et al. (2018); Mihaylov et al. (2018); Zellers et al. (2019); Clark et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Talmor et al. (2019); Bisk et al. (2019); Sap et al. (2019); Hendrycks et al. (2021a); Cobbe et al. (2021); Chen et al. (2021); Austin et al. (2021); Parrish et al. (2022); Zhong et al. (2023); Srivastava et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-03-26 19835 -1355.60 60.09
@@
320 google/gemma-2b Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Zellers et al. (2019); Bisk et al. (2019); Sap et al. (2019); Clark et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Talmor et al. (2019); Gehman et al. (2020); Austin et al. (2021); Hendrycks et al. (2021a); Cobbe et al. (2021); Chen et al. (2021); Dhamala et al. (2021); Parrish et al. (2022); Lin et al. (2022a); Hartvigsen et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 2B 2024-02-08 256798 -581.35 46.51
@@
321 google/gemma-2b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Zellers et al. (2019); Bisk et al. (2019); Sap et al. (2019); Clark et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Talmor et al. (2019); Gehman et al. (2020); Austin et al. (2021); Hendrycks et al. (2021a); Cobbe et al. (2021); Chen et al. (2021); Dhamala et al. (2021); Parrish et al. (2022); Lin et al. (2022a); Hartvigsen et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 2B 2024-02-08 103851 -886.70 42.75
@@
322 google/gemma-7b Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Zellers et al. (2019); Bisk et al. (2019); Sap et al. (2019); Clark et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Talmor et al. (2019); Gehman et al. (2020); Austin et al. (2021); Hendrycks et al. (2021a); Cobbe et al. (2021); Chen et al. (2021); Dhamala et al. (2021); Parrish et al. (2022); Lin et al. (2022a); Hartvigsen et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Dettmers et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-02-08 72047 -539.12 63.75
@@
323 google/gemma-7b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Zellers et al. (2019); Bisk et al. (2019); Sap et al. (2019); Clark et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Talmor et al. (2019); Gehman et al. (2020); Austin et al. (2021); Hendrycks et al. (2021a); Cobbe et al. (2021); Chen et al. (2021); Dhamala et al. (2021); Parrish et al. (2022); Lin et al. (2022a); Hartvigsen et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-02-13 66776 -1327.43 53.56
324 google/recurrentgemma-2b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Zellers et al. (2019); Bisk et al. (2019); Sap et al. (2019); Clark et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Talmor et al. (2019); Gehman et al. (2020); Hendrycks et al. (2021b); Austin et al. (2021); Hendrycks et al. (2021a); Cobbe et al. (2021); Chen et al. (2021); Dhamala et al. (2021); Parrish et al. (2022); Lin et al. (2022a); Hartvigsen et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); De et al. (2024); Griffin Team et al. (2024) recurrent_gemma 2B 2024-04-08 4046 -1033.09 40.86
@@
325 gradientai/Llama-3-8B-Instruct-Gradient-1048k Ding et al. (2023); Peng et al. (2023b); Pekelis et al. (2024); AI@Meta (2024); Liu et al. (2025) llama-3 8B 2024-04-29 6854 -564.98 59.84
326 gradientai/Llama-3-8B-Instruct-262k Ding et al. (2023); Peng et al. (2023b); Pekelis et al. (2024); AI@Meta (2024); Liu et al. (2025) llama-3 8B 2024-04-25 12536 -555.17 60.26
@@
366 HuggingFaceH4/zephyr-7b-alpha Ding et al. (2023); Tunstall et al. (2023); Rafailov et al. (2024); Cui et al. (2024) mistral 7B 2023-10-09 12911 -559.73 59.50
367 HuggingFaceH4/zephyr-7b-beta Ding et al. (2023); Tunstall et al. (2023); Rafailov et al. (2024); Cui et al. (2024) mistral 7B 2023-10-26 294019 -570.40 59.08
@@
375 hywu/Camelidae-8x13B Houlsby et al. (2019); Dettmers et al. (2023); Komatsuzaki et al. (2023); Wu et al. (2024) camelidae 13B 2024-01-10 1886 -535.46 59.40
376 hywu/Camelidae-8x7B Houlsby et al. (2019); Dettmers et al. (2023); Komatsuzaki et al. (2023); Wu et al. (2024) camelidae 7B 2024-01-10 1901 -565.23 54.47
@@
387 IDEA-CCNL/Ziya-LLaMA-13B-Pretrain-v1 IDEA-CCNL (2021); Yang et al. (2022); Zhang et al. (2022a) llama-1 13B 2023-06-01 1171 -1397.30 29.96
388 IDEA-CCNL/Ziya-LLaMA-13B-v1 IDEA-CCNL (2021); Yang et al. (2022); Zhang et al. (2022a) llama-1 13B 2023-05-16 1244 -1397.34 29.82
@@
423 jphme/orca_mini_v2_ger_7b Su et al. (2023b); Mathur (2023a); Conover et al. (2023); Touvron et al. (2023a); Harries (2023); Xu et al. (2023a); Taori et al. (2023) llama-1 7B 2023-07-04 1181 -603.44 47.65
@@
566 meta-math/MetaMath-Llemma-7B Yu et al. (2024b); Azerbayev et al. (2024) llama 7B 2023-11-19 1322 -612.21 53.19
@@
601 mosaicml/mpt-7b Henry et al. (2020); Shoeybi et al. (2020); Dao et al. (2022); Press et al. (2022); Touvron et al. (2023a); Chen et al. (2023a); MosaicML NLP Team (2023b) mpt 7B 2023-05-05 31480 -548.53 44.28
602 mosaicml/mpt-7b-chat Henry et al. (2020); Press et al. (2022); Dao et al. (2022); MosaicML NLP Team (2023b) mpt 7B 2023-05-04 88069 -647.13 45.39
603 mosaicml/mpt-7b-instruct Henry et al. (2020); Press et al. (2022); Dao et al. (2022); MosaicML NLP Team (2023b) mpt 7B 2023-05-05 8599 -569.62 44.83
604 mosaicml/mpt-7b-storywriter Press et al. (2022); Dao et al. (2022); Chen et al. (2023a); MosaicML NLP Team (2023b) mpt 7B 2023-05-04 1920 -650.05 39.31
605 mosaicml/mpt-7b-8k Henry et al. (2020); Shoeybi et al. (2020); Dao et al. (2022); Press et al. (2022); Touvron et al. (2023a); Chen et al. (2023a); MosaicML NLP Team (2023b) mpt 7B 2023-06-30 2106 -546.57 47.24
606 mosaicml/mpt-7b-8k-chat Henry et al. (2020); Press et al. (2022); Dao et al. (2022); MosaicML NLP Team (2023a) mpt 7B 2023-06-22 1304 -586.09 47.78
@@
651 openaccess-ai-collective/jackalope-7b Mukherjee et al. (2023); Longpre et al. (2023); Lian et al. (2023b) mistral 7B 2023-10-07 1198 -550.33 61.16
@@
687 OpenModels4all/gemma-1.1-7b-it Joshi et al. (2017); Zhao et al. (2018); Mihaylov et al. (2018); Zellers et al. (2019); Clark et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Talmor et al. (2019); Bisk et al. (2019); Sap et al. (2019); Hendrycks et al. (2021a); Cobbe et al. (2021); Chen et al. (2021); Austin et al. (2021); Parrish et al. (2022); Zhong et al. (2023); Srivastava et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-04-06 5128 -1355.60 59.78
@@
692 Open-Orca/LlongOrca-13B-16k Wang et al. (2023a); Mukherjee et al. (2023); Dale et al. (2023); Touvron et al. (2023b); Longpre et al. (2023) llama-2 13B 2023-08-16 1217 -551.07 56.59
693 Open-Orca/LlongOrca-7B-16k Wang et al. (2023a); Lian et al. (2023c); Mukherjee et al. (2023); Touvron et al. (2023b); Longpre et al. (2023) llama-2 7B 2023-08-05 1223 -583.06 53.02
@@
695 Open-Orca/Mistral-7B-SlimOrca Mukherjee et al. (2023); Longpre et al. (2023); Lian et al. (2023f, e) mistral 7B 2023-10-08 3696 -569.42 60.37
696 Open-Orca/OpenOrcaxOpenChat-Preview2-13B Wang et al. (2023a); Mukherjee et al. (2023); Touvron et al. (2023b); Longpre et al. (2023); Wang et al. (2023b) llama-2 13B 2023-07-31 2172 -547.85 56.70
697 Open-Orca/OpenOrca-Platypus2-13B Hu et al. (2022); Wang et al. (2023a); Lee et al. (2023); Mukherjee et al. (2023); Touvron et al. (2023b); Longpre et al. (2023); Wang et al. (2023b); Lee et al. (2024) llama 13B 2023-08-11 6853 -557.72 57.28
698 Open-Orca/OpenOrca-Preview1-13B Mukherjee et al. (2023); Longpre et al. (2023); Lian et al. (2023a); Touvron et al. (2023a) llama-1 13B 2023-07-12 1240 -679.87 51.38
@@
742 rinna/youri-7b Andonian et al. (2021); Zhao et al. (2023a); Touvron et al. (2023b); Sawada et al. (2024) llama-2 7B 2023-10-30 2512 -545.03 47.11
@@
766 SeaLLMs/SeaLLM-7B-v2 Zhang et al. (2023); Zheng et al. (2023); Kojima et al. (2023); Nguyen et al. (2024) mistral 7B 2024-01-29 6294 -548.56 67.57